OpenStudy (anonymous):

write the equation of the line with y-intercept of 3 and slope 6

OpenStudy (accessdenied):

Use slope intercept form: y = mx + b. m = slope, b = y-intercept value

OpenStudy (anonymous):

can you help me work through it

OpenStudy (accessdenied):

Well, what's your slope in the given information? We'll insert this into the equation "y=mx+b" as m.

OpenStudy (anonymous):

well the 3 would go into y= 6 +3

OpenStudy (accessdenied):

That is (almost) correct, you just keep the x in the equation next to 6. y = 6x + 3

OpenStudy (anonymous):

y= 6x=3

OpenStudy (anonymous):

how bout this.. show a circle mapping of the relation {(x,yl y = lxl = -3, -1,1} and tell whether or not it's a function